This sentence no verb. Check forspaces, and inclusion of proper texts. This is a line of text **Comment: 1 line(s) removed** another line of text **Comment: 3 line(s) removed** last line of text Let’s test paragraphs: For starters. On the other hand: This is one is present. As is this one. And that’s it.